### v3

# LAuReL: Learned Augmented Residual Layer

Abstract
One of the core pillars of efficient deep learning methods is architectural improvements, such as residual/skip connections, which have led to significantly better model convergence and quality. Since their introduction, residual connections have become ubiquitous not only in convolutional neural networks but also in transformer-based architectures, the backbone of LLMs.
In this paper, we introduce the Learned Augmented Residual Layer (LAuReL)—a novel generalization of the canonical residual connection—designed to serve as an in-situ replacement while outperforming it in both model quality and footprint metrics. Our experiments show that LAuReL can enhance model quality for both vision and language models while adding fewer parameters and incurring less latency and memory overhead than naively increasing parameter count.
For example, on the ImageNet-1K task, LAuReL achieves the same model quality improvements as naively adding an extra layer while using 2.6×2.6$\times$2.6 × fewer parameters. Similarly, when pre-training 1B and 4B parameter LLMs, LAuReL improves performance on a variety of challenging downstream evaluation tasks by 2.54% to 20.05%, while adding only 0.012% and 0.1% additional parameters, respectively.
Machine Learning, ICML
1 Introduction
Model efficiency is of critical importance in the age of extremely large language and vision models. Even if a given model has impressive quality, its footprint metrics such as train-time compute required, inference latency, resident memory size, peak memory consumption, etc. dictate if it can be experimented with and/or deployed in real-world settings. A large and/or slow model, regardless of its performance on eval tasks, might be infeasible to train and serve, and might not be fit for applications where the user expects responsiveness.
LLMs such as Gemini 1.5 Flash (Gemini-Team et al., 2024), DeepSeek V3 (DeepSeek-A.I. et al., 2024) have been explicitly designed with these efficiencies in mind and consistently outperform much larger models that preceded them.
Consequently, improving the Pareto-frontier of model quality vs footprint, via efficient deep learning methods has been an area of active research in the past few years. Areas of interests span from algorithmic techniques (Menghani, 2023), to efficient hardware (Sze et al., 2017), to identifying best practices around model efficiency (Dehghani et al., 2022), etc.
One of the core pillars of efficient deep learning methods are architectural improvements such as the residual/skip connection, which had led to significantly better model convergence and quality (He et al., ). Since its introduction, the residual connection has become ubiquitous in not just convolutional neural networks but also transformer-based architectures (Vaswani et al., 2017), the backbone of LLMs of today.
In this paper we introduce learned augmented residual layer, LAuReL, which generalizes the canonical residual connection. Recall that deep-learning models with residual connections have a ‘block’ structure, with many blocks chained together between the input and final output; these could be convolution/identity blocks within a ResNet, a transformer block in a transformer encoder/decoder, etc. Within a block, a typical residual connection is given by:
$x_{i+1}=f(x_{i})+x_{i}.$ (1)
Here, $f(\cdot)$ can be any non-linear function such as attention, MLP, multiple non-linear layers, etc., $x_{i}$ is the input to the said non-linear function, and $x_{i+1}$ is the combined output of the non-linear function and the residual component. Refer to Figure 1 for an illustration. To simplify exposition, we ignore pre-processing functions such as layer norm, which can be folded into $f(\cdot)$ without loss of generality.
Figure 1: A standard residual connection. We assume the model to be divided into logical ‘blocks’, which is true for most modern architectures including transformers. The residual connection combines the output of a non-linear function $f$ and the input to the said non-linear function. Here, $f$ can be attention, MLP, or any other combination of non-linear layers.
2 Learned Augmented Residual Layer
In this section we describe the main idea behind LAuReL. In its most general form, we reformulate the residual connection to be the following:
$x_{i+1}=\alpha\cdot f(x_{i})+g(x_{i},x_{i-1},\ldots,x_{0}).$ (2)
Here $\alpha$ is a learned scalar parameter, and $g(\cdot)$ is a learned linear function with $x_{i},x_{i-1},\ldots,x_{0}$ as inputs, where $x_{j}$ is the output of the $j$th residual connection. The intuition behind LAuReL is that one can learn a richer set of (linear) functions than just using $x_{i}$ as the residual component. One motivation behind seeking these richer linear functions is the concept of a “residual stream” (Elhage et al., 2021), where the residual connection is considered to be part of a stream of information that passes through each layer without being exposed to any non-linearities. This allows the learning process to focus on the non-linear components better.
Each layer / operation can read from, and subsequently write to this residual stream based on what it read. Given that the residual connection has been shown to be important for model quality and convergence, we designed LAuReL to operate on this residual stream in a learned fashion, while being light-weight in terms of the model size and latency changes.
Figure 2: An illustration of the LAuReL framework; see (2). LAuReL can be used to replace the regular residual connection in Figure 1. Again, $f$ can be any non-linear function such as attention, MLPs, and groups of multiple non-linear layers.
In this paper we study three specific versions of the LAuReL framework; although as described in (2), the framework can be generalized beyond these versions.
2.1 Residual Weights Version (LAuReL-RW)
In this version, we keep $\alpha$ learnable and set $g(x_{i},\ldots,x_{0})=\beta x_{i}$. Therefore, (2) can be rewritten as:
$x_{i+1}=\alpha f(x_{i})+\beta x_{i}.$
Notice that this version assigns learnable weights to the $f(x_{i})$ and $x_{i}$ from (1). In practice, we found that we cannot let $\alpha$ and $\beta$ grow unbounded, and using a normalization function such as softmax or sigmoid helps. Clearly, this version will add only two new parameters per LAuReL layer. If necessary, we can always replace these two parameters by a single learnable parameter and use a function like sigmoid to define $\alpha,\beta$ in terms of this single parameter to create a convex combination.
2.2 Low-Rank Version (LAuReL-LR)
In this version, we fix $\alpha=1$, and $g(x_{i})=Wx_{i}$ in (2) to obtain:
$x_{i+1}=f(x_{i})+Wx_{i}.$
Note that, as written, $W$ is a learnable $D\times D$ matrix, where $D$ is the model dimension for transformer-based models, or more generally it is the last dimension of $x_{i}$; hence the $W$ matrix will add $D^{2}$ new parameters (per LAuReL layer) to the model.
In practice, to reduce the number of new parameters added to the model and to help with convergence, we consider a low rank version of $W$. In particular, let $W=A\times B+I$, where $A$ and $B^{T}$ are $D\times r$ matrices and $r\ll D$. Thus, we can rewrite (2) as:
$x_{i+1}=f(x_{i})+BAx_{i}+x_{i}.$ (3)
Here, both $A$ and $B$ matrices are learnable. The number of new parameters is $2rD$.
2.3 Previous Activations Version (LAuReL-PA)
This version is similar to LAuReL-LR, except that we use use $k$ activations from the previous blocks. In particular, we set
$g(x_{i})=(\sum_{j=0}^{k-1}\gamma_{i,j}\cdot h_{i}(x_{i-j}))+x_{i}.$
Where $\gamma_{i,0},\ldots,\gamma_{i,k-1}$ are learned scalar parameters and $h_{i}$ is another linear function. [^1] For simplicity, we fix $\alpha=1$. This allows us to rewrite (2) as:
$x_{i+1}=f(x_{i})+(\sum_{j=0}^{k-1}\gamma_{i,j}\cdot h_{i}(x_{i-j}))+x_{i}.$ (4)
In practice, we replace $h_{i}$ by a low-rank product similar to the LAuReL-LR version, but using the identity function is also an option. When using the low-rank product for $h_{i}$, the number of new parameters is $2rD+k$, where $k$ is the number of previous activations we are accessing.
2.4 Other Derived Variants
Note that, all three proposed versions mentioned above are a combination of scalar and/or low-rank products on top of the vanilla residual connection in (1). This makes LAuReL especially light-weight in terms of its impact on model size and latency. We discuss the efficiency of LAuReL variants in more detail in Section 4.
That being said, our framework is generic enough to allow combinations of the above variants, as well as new variants. For instance, one straight-forward combination is LAuReL-RW+LR, where the residual weights from LAuReL-RW can be applied along with the LAuReL-LR variant to rewrite (3) as follows:
$x_{i+1}=\alpha f(x_{i})+\beta(BAx_{i}+x_{i}).$ (5)
Similarly, LAuReL-PA as defined in (4) can be combined with LAuReL-RW as follows:
$x_{i+1}=\alpha f(x_{i})+\beta((\sum_{j=0}^{k-1}\gamma_{i,j}\cdot h_{i}(x_{i-j}))+x_{i}).$
When using a low-rank product for $h_{i}$, we can create the LAuReL-RW+LR+PA variant as follows:
$x_{i+1}=\alpha f(x_{i})+\beta((\sum_{j=0}^{k-1}\gamma_{i,j}\cdot A_{i,j}B_{i,j}(x_{i-j}))+x_{i}).$ (6)
Yet another variant while slightly relaxing our above formulations would be to treat $\alpha$ and $\beta$ as vectors of length $D$, which is the size of the model dimension in case of transformer-like models. In other cases it is the size of the last dimension of the input $x_{i}$. In this variant we would learn fine-grained per-dimension weights when mixing $x_{i}$ and $f(x_{i})$ at the cost of $2D$ parameters per LAuReL-LR layer, instead of two per parameters per layer.
To summarize, LAuReL is inherently flexible and provides many possible cheap learnable augmentations on top of the vanilla residual connection. In the following section we demonstrate that these combinations are efficient and effective at improving the model quality of common vision and language models, while having a minimal impact on the number of parameters and latency.
3 Experiments
We experiment with LAuReL in two domains, namely, vision and language. For the first case, our goal is to improve the image classification accuracy of the ResNet-50 model on the ImageNet-1K dataset (Deng et al., 2009). For the second case our goal is to improve the performance of two different large language models (LLMs) of size 1B and 4B params respectively, evaluated after the pre-training stage, on common benchmarks.
The underlying motivation behind these experiments is not necessarily to improve on the SOTA results, but to show how LAuReL can be easily integrated on top of common model architectures with residual/skip connections in order to achieve a better model quality and footprint trade off.
3.1 ResNet-50 on ImageNet-1K
In this setup we train a standard ResNet-50 model on the ImageNet 1K dataset (Deng et al., 2009) using 16 Google Cloud TPUv5e chips over one epoch with data-augmentation turned on. In order to obtain a strong baseline, we fine-tuned the model learning rate schedule and picked a schedule that maximized the average of the best accuracy@1 values over 5 trials (which we simply refer to as accuracy in this subsection). The baseline model that we obtained achieves an accuracy of $74.95\pm 0.016\%$.
In addition, we also find that if we simply add another layer to the ResNet-50 model (i.e., naive scaling), we can increase the model’s accuracy by $0.25\%$ to reach $75.20\%$, while adding $4.37\%$ new parameters. With that in context, applying LAuReL on the model leads to better results (see Table 1).
If we only use the LAuReL-RW version, we get an improvement of $0.15\%$ on average with only $0.003\%$ extra parameters, which is essentially negligible. When we try the LAuReL-RW+LR version from (5) with $r=16$, we achieve an accuracy of $75.20\%$ while adding only $1.68\%$ extra parameters; this matches the performance of the baseline with an extra layer, while using 2.6×2.6$\times$2.6 × fewer extra parameters. Additionally, when we use the combined LAuReL-RW+LR+PA version from (6) we improve the accuracy to $75.25\%$ while still using 1.82×1.82$\times$1.82 × fewer extra parameters than the baseline with one extra layer, demonstrating that LAuReL is superior to naively scaling the model. Notably even though we make fundamental changes to the residual connection we did not find any training instabilities when using LAuReL.
Table 1: Applying LAuReL on a ResNet-50 trained on the ImageNet 1K classification dataset. LAuReL-RW provides a significant boost with negligible extra parameters. LAuReL-RW+LR and LAuReL-RW+LR+PA meet and beat the naive scaling baseline while using 2.6×2.6$\times$2.6 × and 1.82×1.82$\times$1.82 × fewer parameters.
Model Avg. Best Params Added
Accuracy@1 vs Baseline
(%), 5 trials (%)
Baseline $74.95\pm 0.01$ $0.00$
Baseline + 1 Layer $75.20\pm 0.12$ $4.37$
(Naive Scaling)
LAuReL-RW $75.10\pm 0.10$ $0.003$
LAuReL-RW+LR $\bf 75.20\pm 0.07$ $\bf 1.68$
LAuReL-RW+LR+PA $\bf{75.25\pm 0.09}$ $\bf 2.40$
3.2 Large-Scale LLM Pre-training
In this setup, our goal was to test the performance of LAuReL when applied on top of strong LLMs. During the course of our work we evaluated LAuReL on two separate LLMs, which we pre-trained from scratch. The first LLM (LLM-1) was a 1B parameter model pre-trained with a data-mixture consisting of text tokens only. The second LLM (LLM-2) was a 4B parameter model which was pre-trained with a multi-modal and multi-lingual data mixture. Both LLMs were trained with $\Theta(0.5T)$ tokens.
What we varied across the two LLMs was to allow for different budgets for increasing model footprint (params, latency, memory, etc.) when applying LAuReL. For LLM-1 we allow a very small increase in these metrics ($\sim$ 0.01% extra params, and nearly no latency increase). For LLM-2 we allow a slightly lenient, yet very modest increase in params ($\sim$ 0.1% extra params), and latency (1% increase).
Given the scale of LLMs today, both the budgets would be considered negligible. For instance, a $0.1\%$ increase in parameters for a 4B model corresponds to 4M more parameters. As demonstrated in the ResNet / ImageNet experiments in Section 3.1, LAuReL outperforms naive scaling, but we further detail the comparison with naive scaling in Section 4.4.
Our objective behind testing LAuReL in these conditions was to demonstrate the efficacy of the method and ensure that LAuReL scales well across different LLM setups in the wild.
3.2.1 LLM-1: 1B params, $\Theta(0.5\textrm{T})$ tokens, very low extra params/latency/memory budget
For our first baseline, we chose a 1B parameter decoder-only model based on the Transformer architecture. We pre-trained both the baseline, and our experiment with LAuReL, from scratch; we use the LAuReL-RW and LAuReL-LR versions (with $r=4$). Both the models were trained using 256 Google Cloud TPU v5e chips for approximately two weeks each, using a pre-training mixture consisting of only text data which included webpages, books, code, translations, and other text data.
It is worth noting that the combined LAuReL-RW+LR variant adds only 0.012% more parameters as compared to the baseline model. Since we chose $r=4$, the total number of parameters added by LAuReL-LR is $8ND$ and $2N$ parameters by LAuReL-RW, for a total of $2N(4D+1)$ parameters. Typically $N\in[10,100]$ and $D\in[500,5000]$. For the sake of illustration, assume $N=20$ and $D=1000$, using the LAuReL-RW+LR leads to $160,040$ extra params in a 1B param model. Thus, the number of new parameters is dwarfed by that of the original model. Additionally, we found that the additional latency introduced by LAuReL-RW+LR was within the noise range.
We evaluated both the pre-trained baseline and LAuReL models on a host of common LLM tasks such as Q&A, NLU, Math, Code, etc; see Table 2 for the results. The task type and individual tasks are listed in the first and second columns respectively, and a higher score is better for all the tasks. LAuReL outperforms the baselines on all tasks except on the MBPP dataset where it was neutral. To reiterate, these improvements were achieved with only 0.012% extra parameters and nearly no increase in latency.
Table 2: Evaluation results of LLM-1 as described in Section 3.2.1; a 1B parameter Decoder-only LLM pre-trained from scratch with (a) once with the baseline model architecture, and (b) once using LAuReL on top. We evaluated both the models on a number of common eval benchmarks (higher is better for all task type and task combinations listed below). The LAuReL variant outperforms the baseline on all but one dataset while adding only 0.012% extra parameters.
Task Type Task Baseline LAuReL
Math Math 3.54 3.70
(Hendrycks et al., 2021b) (+4.51%)
GSM8K-CoT 8.34 8.79
(Cobbe et al., 2021) (+5.39%)
General MMLU 25.72 25.89
Reasoning (Hendrycks et al., 2021a) (+0.06%)
Q&A BoolQ 58.07 65.66
(Clark et al., 2019) (+13.08%)
TyDi QA 67.98 72.58
(GoldP) (+6.07%)
(Clark et al., 2020)
Sentence Hellaswag 64.84 65.06
Completion (Zellers et al., 2019) (+0.03%)
Code HumanEval 18.29 18.90
(Chen et al., 2021) (3.33%)
MBPP 27.00 27.00
(Austin et al., 2021)
GSM8K-PAL 10.31 11.37
(Cobbe et al., 2021) (10.28%)
Table 3: Evaluation results of LLM-2 as described in Section 3.2.2; a 4B parameter Decoder-only LLM pre-trained from scratch with (a) once with the baseline model architecture, and (b) once using LAuReL on top. We evaluated both the models on a number of common eval benchmarks (higher is better for all task type and task combinations listed below). The LAuReL variant outperforms the baseline on all but two tasks, while adding only $\sim$ 0.1% extra parameters.
Task Type Task Baseline LAuReL
Math Math 14.70 15.30
(Hendrycks et al., 2021b) (+4.08%)
MGSM 20.0 23.09
(Shi et al., 2023) (+6.07%)
General MMLU 49.85 51.12
Reasoning (Hendrycks et al., 2021a) (2.54%)
Reading Belebele 58.40 63.23
Compre- (Bandarkar et al., 2024) (+8.27%)
hension BookQA 50.36 60.46
(Mihaylov et al., 2018) (+20.05%)
Translation WMT23 68.32 68.24
(Kocmi et al., 2023) (-0.11%)
Multimodal MMMU 32.22 36.33
(Yue et al., 2024) (+12.75%)
Coco-Caption 95.69 99.15
(Lin et al., 2014) (+3.61%)
DocVQA 68.28 68.34
(Mathew et al., 2021) (+0.08%)
TextVQA 60.07 62.64
(Singh et al., 2019) (+4.27%)
3.2.2 LLM-2: 4B params, $\Theta(0.5\textrm{T})$ tokens, low extra params/latency/memory budget
In this second setting, we worked on a bigger 4B parameter decoder-only model with a similar token budget, but trained on a multimodal and multilingual corpus of tokens.
To compensate for a 4×4$\times$4 × larger model, and also the fact that the dataset and eval tasks are harder, we allowing a bigger footprint budget for LAuReL; $\sim 0.1\%$ extra parameters and $\sim 1\%$ extra latency. As mentioned earlier, this is still a negligible increase in model params and latency.
To match these budgets, we set $r=64$, which provides a favorable trade-off of providing more capacity to the low-rank matrices as specified in the LAuReL-RW+LR formulation in (6), while also meeting the above-mentioned budgets.
In this setting, both the baseline and the LAuReL experiment were trained using 1024 Google Cloud TPU v4 chips for slightly more than two days each. See Table 3 for the evaluation results of both the baseline model and the model with LAuReL. The task type and individual tasks are listed in the first and second columns respectively, and a higher score is better for all the tasks. LLM-2 had a sophisticated suite of evaluation tasks, including math, general reasoning, reading comprehension, translation, and multimodal tasks. All of the listed eval tasks are used by leading LLMs to evaluate model quality, further solidifying our confidence in the model evaluation setup.
To start off, two of the LLM-2 evaluation tasks, Math, and MMLU were common with LLM-1 (refer to Table 2 for LLM-1 results). It can be seen that LLM-2 is much stronger than LLM-1 on both the common evals; 4.16×4.16$\times$4.16 × better on the Math task, and 1.93×1.93$\times$1.93 × better on the MMLU task. We attribute this to LLM-2 being 4×4$\times$4 × larger, and having a more sophisticated pre-training mixture. LAuReL using the LAuReL-RW+LR variant improves on both Math and MMLU, which is pleasantly surprising given how powerful is LLM-2. For other tasks LAuReL improves significantly over the baseline as well, except for WMT23 and DocVQA, where it was neutral.
In terms of costs, the model adds $\sim 0.1\%$ more parameters. This is still a very reasonable approximately 4M parameters on top of a 4B parameter model. In terms of latency, we tested on both server-side (Google CloudTPU) for cloud serving, and a leading smartphone for on-device inference. In both the benchmarks, we measure nearly 1-2% increase in latency for prefill and generation. There was no human perceptible difference in terms of time-to-first-token.
We did not try the LAuReL-PA version for the above LLM experiments, as the LLM training was expensive. Additionally the LAuReL-PA variant requires slightly more memory to keep previous block activations accessible. This also meant changing sharding, rematerialization, and other settings in the LLM codebases we worked on. However, we expect the LAuReL-PA results from the ResNet experiments to also hold in this case as well.
3.3 LAuReL-LR: Rank vs Accuracy
We note that for the LAuReL-LR version on the ResNet-50/ImageNet combination, there is a pattern in terms of the best accuracy achieved with different values of $r$. In the combined LAuReL-RW+LR version we experimented with different values of $r$, and computed the average of the best accuracy@1 achieved over 5 trials; see Figure 3. From Table 1, with the LAuReL-RW version alone we already achieve an average best accuracy@1 of 75.10%, therefore for the combined LAuReL-RW+LR version we would like to see the accuracy exceeding that.
We observe that when $r$ is small ($r\in\{4,8\}$), there is not a significant improvement over the baseline LAuReL-RW experiment. This could be because a very small $r$ acts as an information bottleneck in the low-rank product in (3). As $r$ increases, the accuracy reaches the maximum for $r=16,32$; beyond this, the accuracy seems to drop though still higher than the LAuReL-RW baseline.
We believe this unimodal phenomenon could be due to the number of parameters added to the model, which increases linearly in $r$, this would also require appropriate tuning of hyper-parameters such as the learning rate as well as the regularization penalty.
Another possible cause could be the initialization scheme for the low-rank matrices $A$ and $B$. In our experiments, we set $A$ to be initialized either using Xavier initialization for ResNet, or using an orthogonal initialization for the LLM experiments. $B$ was always initialized to zero. This is similar to the scheme used in Hu et al. (2022). We found that initialization made a significant difference in the performance of the LAuReL-LR variant, and we posit that further work studying and improving the initialization scheme of the low-rank matrices could lead to better performance of the variant.
Figure 3: Trend of accuracy as $r$ is varied in the LAuReL-RW+LR variant.
4 Efficiency
Large models (LLMs and beyond) have been scaled both in the number of parameters as well as the number of tokens to achieve better model quality that scaling laws (Hoffmann et al., 2022) promise. However, this is directly in conflict with keeping training and inference costs reasonably low (we describe these costs shortly). These competing forces have resulted in the emergence of models such as Gemini Flash and Nano (Gemini-Team et al., 2024), DeepSeek (DeepSeek-A.I. et al., 2024), etc., which have attractive cost versus quality trade-offs.
To achieve these favorable trade-offs, the above-mentioned models incorporate training techniques and architectural changes that can help improve quality while keeping training and inference cheap enough. Therefore, any proposed model efficiency technique including LAuReL should demonstrate not just improvement in quality, but also training and inference efficiency.
4.1 Efficiency Metrics
We now define some of the metrics on which we can measure model training and inference efficiency, before describing how LAuReL scores on them.
4.1.1 Number of Parameters
This is the most commonly quoted metric that is designed to capture the cost of training and serving the model. When the model in question is an MoE, there are two numbers quoted: the total number of parameters and the number of activated parameters. Note that a larger number of parameters implies greater forward and backward pass costs.
4.1.2 Latency (Training and Inference)
For training, a reasonable metric to consider would be the number of steps taken per second. Similarly, an important metric would be inference latency when deploying the model. For LLMs this can be broken down into the latency metrics in the warm-up stage (e.g., time-to-first-token, which can be further refined to prefill latency), and the generation stage (output tokens/second). For LLMs and other interactive models, both the time-to-first token and output tokens/second are important for a good user experience.
4.1.3 Memory
Peak memory used during training is another key metric that is tracked to ensure that accelerators can fit in the model graph and have enough free memory available for forward and backward passes. Inefficiencies can be compensated by strategies like rematerialization (Kumar et al., 2019), which can help reduce peak memory usage but at a cost of recomputing some activations. Similarly, peak memory used during inference is also an important metric to track.
4.2 Efficiency Analysis
LAuReL variants are designed with the above efficiency metrics in mind, and in this section we study the performance of the variants on these metrics.
Table 4: Analysis of extra params, memory, and latency incurred for each LAuReL variant, per instantiation, except the LAuReL-PA case where at any time no more than $k$ previous activations are in-memory, incurring a total $\Theta(kD)$ extra memory cost. Also note that the latency cost of LAuReL-LR can be tighter than $O(rD^{2})$, depending upon the exact matrix multiplication algorithm used. The latency complexities ignore the batch and sequence dimensions without loss of generality.
LAuReL Params Memory Latency
variant
[…]
LAuReL-LR $2rD$ $\Theta(2rD)$ $O(rD^{2})$
$D$ denotes model dim,
$r<<D$)
LAuReL-PA $k$ $\Theta(kD)$ $O(kD)$
(k denotes number of
of prev. activations)
LAuReL-LR+PA $2krD$ $\Theta(2krD$ $O(krD^{2}$
[…]
Table 4 lists the costs associated with using each LAuReL layer. We assume that we are working with a vector input of size $D$ (the last dimension of the input). This should not have any impact on the calculations or the generalization.
Note that the listed costs are per-LAuReL instantiation. Hence, if there are many transformer layers and if a particular LAuReL variant is used in each layer, then the relevant costs accumulate across the layers. We now examine the costs of the LAuReL variants.
LAuReL-RW is the cheapest with one or two scalars as parameters, a constant memory cost (the constant depending on the data type: fp32, fp16 / bf16, unit8 / int8, etc.), and a constant latency cost since the scalar multiplication might be optimized by the compiler.
The LAuReL-LR variant has $2rD$ parameters per instantiation ($rD$ for each of the two low-rank matrices, $A$ and $B$), and uses $\Theta(rD)$ memory. The latency is upper bounded by $O(rD^{2})$, since the input is a vector of dimension $D$ (ignoring the batch and sequence dimensions), and there are two matrix multiplications.
Finally, the LAuReL-PA variant uses $k$ previous activations. Assuming $h_{i}$ is an identity function, it uses $k$ extra parameters, and uses $\Theta(kD)$ extra memory for the previous activations. Since it requires $k$ extra vector additions, this introduces an additional latency of $O(kD)$.
4.3 Practical Recommendations
Given the experiments and the analysis of individual variants, LAuReL-RW would be the obvious first candidate to try. The LAuReL-RW+LR offers further improvement in model quality, and seems to provide the best trade-off in terms of model quality improvements and the overhead of parameters, latency, and memory. We validated this in ResNet and LLM experiments.
The LAuReL-RW+LR+PA variant also leads to improvements over LAuReL-RW+LR as the ResNet experiments show. The main caveat is the additional $\Theta(kD)$ memory, which might be undesirable. However, this can be mitigated to some extent by choosing a small value of $k$ or by changing the model sharding/rematerialization methods.
4.4 Comparison with Naive Model Scaling
One of the advantages of LAuReL to highlight is that it is competitive against naive scaling. Concretely, given a modest additional budget for parameters, latency, memory, allocating that budget to LAuReL (or its variants) is likely to produce more gains than naively using it in an existing model architecture.
We observed this in the ResNet experiments (Section 3.1), where LAuReL variants match the performance of naive scaling while adding 2.6$\times$ fewer parameters, thus showing the Pareto-efficiency of LAuReL (vs naive scaling).
However, in the LLM experiments in Section 3.2, the gains achieved using LAuReL variants are at the cost of $\sim$ 0.01–0.1% more parameters. A naive way to use this ’additional’ budget would be, say, increase the model dimension ($D$). However, this is may not be feasible due to hardware limitations and memory alignment issues. Another way to use this ’additional’ budget would be to increase the vocabulary size. However, the gains would be limited since this will include more tokens only from the tail of the distribution and may not contribute much to the model quality. Interestingly, Karpathy (2023) reported that the NanoGPT model (Karpathy, 2022) pre-training was sped up by $\sim$ 25% when the vocabulary size was made divisible by 64.
To summarize, not only did we demonstrate that LAuReL outperforms naive scaling, but also in transformer-style models, naively distributing the relatively small additional budget of LAuReL parameters might actually hurt performance and would require non-trivial systems engineering.
5 Related Work
Since our larger goal is to improve the training and inference efficiency of deep learning models, we briefly discuss some research directions aimed at improving model efficiency.
Architectural Changes: Our work is inspired by recent model architecture improvements such as LoRA (Hu et al., 2022) and AltUp (Baykal et al., 2023) amongst others. However, they are not directly relevant to LAuReL. Indeed, LoRA is designed to efficiently fine-tune large pre-trained models and it works directly on the model weight matrices level by introducing low-rank ‘adapter’ weights that are learned during the fine-tuning stage, while other model weights are held constant. In contrast, LAuReL works at the residual connection level, which likely spans multiple weight matrices involved in the function $f$; furthermore, it is applied during the pre-training stage.
AltUp (Baykal et al., 2023) is designed to replicate the quality improvements of a model with a large model dimension, without having to pay the additional cost. It operates at the transformer-block level, constructing parallel ‘lightweight’ transformer blocks to approximate the model dimension scaling effect. In contrast, LAuReL operates at the residual connection level and does not aim to replicate the dimension scaling effect.
Interestingly, LAuReL can be applied in conjunction with both LoRA (during fine-tuning) and AltUp (during pre-training and fine-tuning).
He & Hofmann (2024) proposes several changes to transformer blocks to help improve model convergence; however these proposals are limited to transformer blocks.
Compression Techniques: Model compression techniques (Buciluǎ et al., 2006) such as quantization (Krishnamoorthi, 2018; Jacob et al., 2018), including ternary networks (Li et al., 2016; Ma et al., 2024) are commonly used to reduce model size and inference latency. Similarly, pruning and model sparsity (Gale et al., 2019; Liu et al., 2019) techniques have also been explored and implemented in hardware.
Learning Techniques: Distillation (Hinton et al., 2014) is a popular technique for improving a smaller (student) model quality using “soft-labels” from a larger, impractical (teacher) model (Sanh et al., 2019). Some distillation variants propose learning intermediate representations as well (Zagoruyko & Komodakis, 2016; Kim et al., 2023). Other techniques include works like Stacking (Reddi et al., 2023) and RaPTr (Panigrahi et al., 2024), which progressively grow and train the network to achieve an improved model quality while reducing model training time.
6 Conclusion
In this paper we introduce the LAuReL framework, which is a novel architectural change and a generalization of the residual / skip connection aimed at improving the model quality without significantly increasing the model size or latency. We study three versions (LAuReL-RW, LAuReL-LR, LAuReL-PA) that can be mixed-and-matched together, as we show in our experiments.
Through experiments, we demonstrate the efficacy of replacing the conventional residual connection with LAuReL on both vision and language tasks, while also providing evidence for its advantages over naive model scaling methods. For future work, we would like to try to further improve LAuReL with new variants with better quality v/s model footprint trade-offs.